# Safe-LLaVA: A Privacy-Preserving Vision-Language Dataset and Benchmark for Biometric Safety

### v1

@@ 5 Discussion
In this work, we addressed the challenge of biometric privacy in Vision-Language Models (VLMs) through two core contributions: (1) constructing a privacy-preserving dataset, and (2) introducing a benchmark for privacy-aware evaluation. First, we developed the Safe-LLaVA dataset by systematically removing biometric attributes such as eye color, gender, age, race, and body type—while preserving semantic content. Models trained on Safe-LLaVA significantly reduced biometric leakage without compromising general performance, demonstrating the effectiveness of proactive dataset cleaning beyond existing memorization-focused approaches. Second, we proposed PRISM, the first benchmark explicitly designed to assess biometric privacy in VLMs. PRISM evaluates both refusal behavior on direct biometric queries and implicit leakage in open-ended responses. Our experiments show that Safe-LLaVA-trained models achieve higher refusal accuracy and implicit leakage protection, validating the effectiveness of our Safe-LLaVA dataset.
Supplementary: Safe-LLaVA: A Privacy-Preserving Vision-Language Dataset and Benchmark for Biometric Safety
We organize the supplementary material as follows:
@@ Appendix B Implementation Details
Safe-LLaVA (0.5B) shares the same model architecture and training configuration as LLaVA-OneVision (0.5B)Li et al. (2024), and Safe-LLaVA (7B) is identical in architecture and setup to LLaVA-v1.5 (7B)Liu et al. (2024b). Both Safe-LLaVA (0.5B) and Safe-LLaVA (7B) are trained on the proposed Safe-LLaVA dataset using the exact same model settings. The only difference between baseline LLaVA-v1.5 (7B) and Safe-LLaVA (7B) lies in the training data: Safe-LLaVA models are trained on privacy-filtered corpora in which explicit and implicit biometric attributes have been removed.
Appendix C PRISM-Large Results
Table 4 demonstrates refusal accuracy on PRISM-Large across five biometric attributes, evaluated separately under soft and hard prompts. Across both evaluators (GPT-4 and Gemini-2.0-flash), baseline models such as Gemma, Qwen2.5-VL, and LLaVA variants exhibit poor refusal rates, often close to zero, with moderate improvements in larger-scale variants (e.g., LLaVA-v1.5 7B). In contrast, Safe-LLaVA achieves near-perfect or perfect refusal accuracy in almost all categories, demonstrating robustness under both prompt types. Notably, Safe-LLaVA (0.5B) consistently attains 100% refusal across attributes, while Safe-LLaVA (7B) also sustains strong performance with minor variations in gender refusal. These results highlight that Safe-LLaVA’s improvements are not only significant but also demonstrate strong generalizability, as the model maintains robust refusal accuracy even under the more diverse and large-scale settings of PRISM-Large.
In Table 5, we present implicit leakage protection scores at the attribute level under open-ended queries. Baseline models show considerable leakage, with average protection scores ranging between . Even stronger baselines (Qwen2.5-VL, LLaVA-v1.5) reveal vulnerability, particularly in sensitive categories such as gender. In stark contrast, Safe-LLaVA substantially outperforms all baselines, reaching average scores above for both 0.5B and 7B variants. The improvements are especially pronounced in gender leakage protection, where Safe-LLaVA surpasses baselines by more than . This demonstrates Safe-LLaVA’s capability to generalize protection effectively across all biometric attributes, substantially reducing unintended information disclosure.
| Evaluator(Soft) | Model(Param.) | ACC | ACC | ACC | ACC | ACC |
| GPT-4 | Gemma(4B)Team et al. (2024) | 0.1 | 0 | 0 | 0.36 | 0.72 |
| Qwen2.5-VL(7B)Yang et al. (2024) | 0.05 | 0 | 0.21 | 1.64 | 0.77 | |
| LLaVA-OneVision(0.5B)* Li et al. (2024) | 0.05 | 0 | 0 | 0.41 | 0.05 | |
| LLaVA-v1.5(7B)* Liu et al. (2024b) | 12 | 4.82 | 5.49 | 17.18 | 25.64 | |
[…]
| Gemini-2.0-flash | Gemma(4B)Team et al. (2024) | 0 | 0 | 0 | 0.26 | 4.1 |
| Qwen2.5-VL(7B)Yang et al. (2024) | 5.03 | 0.97 | 1.85 | 7.59 | 30.05 | |
| LLaVA-OneVision(0.5B)* Li et al. (2024) | 0 | 0 | 0 | 0.15 | 0 | |
| LLaVA-v1.5(7B)* Liu et al. (2024b) | 11.69 | 4.72 | 4.1 | 18.51 | 26.62 | |
[…]
| Evaluator(Hard) | Model(Param.) | ACC | ACC | ACC | ACC | ACC |
| GPT-4 | Gemma(4B)Team et al. (2024) | 0.1 | 0 | 0 | 0.36 | 0.26 |
| Qwen2.5-VL(7B)Yang et al. (2024) | 0.97 | 0 | 0.05 | 1.44 | 0.05 | |
| LLaVA-OneVision(0.5B)* Li et al. (2024) | 0.05 | 0 | 0 | 0.36 | 0.1 | |
| LLaVA-v1.5(7B)* Liu et al. (2024b) | 5.38 | 1.9 | 0.62 | 2.41 | 22 | |
[…]
| Gemini-2.0-flash | Gemma(4B)Team et al. (2024) | 0 | 0 | 0.05 | 0.36 | 9.64 |
| Qwen2.5-VL(7B)Yang et al. (2024) | 10.15 | 2.05 | 0.67 | 7.64 | 75.08 | |
| LLaVA-OneVision(0.5B)* Li et al. (2024) | 0 | 0 | 0 | 0.31 | 0 | |
| LLaVA-v1.5(7B)* Liu et al. (2024b) | 10.31 | 1.69 | 0.36 | 5.95 | 30.31 | |
[…]
| Evaluator | Model(Param.) | L | L | L | L | L | Average |
| GPT-4 | Gemma(4B)Team et al. (2024) | 23.37 | 12.58 | 67.54 | 91.32 | 96.58 | 58.28 |
| Qwen2.5-VL(7B)Yang et al. (2024) | 80.44 | 68.19 | 96.0 | 98.63 | 99.44 | 88.54 | |
| LLaVA-OneVision (0.5B)* Li et al. (2024) | 73.15 | 4.8 | 97.09 | 99.03 | 99.3 | 74.67 | |
| LLaVA-v1.5 (7B)* Liu et al. (2024b) | 71.42 | 9.28 | 95.62 | 97.62 | 99.5 | 74.69 | |
[…]
| Gemini-2.0-flash | Gemma(4B)Team et al. (2024) | 22.63 | 34.67 | 61.83 | 86.55 | 95.47 | 60.23 |
| Qwen2.5-VL(7B)Yang et al. (2024) | 79.57 | 77.4 | 94.82 | 93.13 | 98.58 | 88.7 | |
| LLaVA-OneVision (0.5B)* Li et al. (2024) | 73.16 | 23.61 | 98.94 | 97.54 | 99.35 | 78.52 | |
| LLaVA-v1.5 (7B)* Liu et al. (2024b) | 70.65 | 24.82 | 98.36 | 96.21 | 99.42 | 77.89 | |
[…]
| Evaluator | Models (Param.) | Lsentence |
| GPT | Gemma (4B) | 5.06 |
| Qwen2.5 (7B) | 64.51 | |
| LLaVA-OneVision (0.5B)* | 3.01 | |
| LLaVA-v1.5 (7B)* | 7.95 | |
[…]
| Gemini | Gemma (4B) | 5.57 |
| Qwen2.5 (7B) | 60.53 | |
| LLaVA-OneVision (0.5B)* | 19.37 | |
| LLaVA-v1.5 (7B)* | 19.98 | |
[…]
We also report sentence-level implicit leakage protection, where any response that reveals a biometric attribute is counted as incorrect. Table 6 shows that Safe-LLaVA achieves near-perfect sentence-level implicit leakage protection, reaching 98.8 (0.5B) and 96.9 (7B) with GPT evaluation and 95.9 (0.5B) and 93.5 (7B) with Gemini evaluation. These results indicate that Safe-LLaVA does not reveal biometric information in natural language responses, even under open-ended prompts, thereby ensuring strong protection at the sentence level.
Appendix D Qualitative Examples
D.1 Images in PRISM Benchmark
@@
D.2 Qualitative Examples of generated sentences from LLaVA-v1.5(7B) and Safe-LLaVA(7B)
@@
Appendix E Prompts for Safe-LLaVA Dataset Curation
@@
E.1 Prompts for PRISM Benchmark
@@
The full prompt texts used to guide GPT-4o and Gemini evaluators are shown in Figures 14 and 15, which provide step-by-step rules, visual examples, and output formatting constraints.
@@
Appendix F Experiment Statistical Significance
To assess the consistency and reliability of model behavior under the PRISM benchmark, we report the standard deviation of refusal accuracy and implicit leakage protection score across 3 independent runs per model and per prompt. For each prompt, both LLaVA-v1.5(7B) and Safe-LLaVA(7B) generated responses three times, and the resulting outputs were evaluated using two independent evaluators: GPT-4 and Gemini-2.0-flash.
Error Margin:
Inference Time:
We also report average response generation time per prompt. In refusal settings(Table 7), Safe-LLaVA(7B) shows slightly longer inference time compared to LLaVA-v1.5(7B), which is attributed to Safe-LLaVA generating longer refusal sentences (e.g., “I am so sorry, but I cannot provide biometric information”) versus LLaVA’s short one-word answers (e.g., Female”, “White”). In contrast, in open-ended responses(Table 8), Safe-LLaVA(7B) is slightly faster than LLaVA-v1.5(7B). This is likely due to Safe-LLaVA generating privacy-filtered descriptions that skip redundant or speculative biometric details, while still maintaining semantic richness.
These results demonstrate that Safe-LLaVA not only improves privacy protection with high consistency, but does so with minimal computational overhead.
| Evaluator(Soft) | Model(Param.) | ACC | ACC | ACC | ACC | ACC | Time(sec) |
| GPT-4 | LLaVA-v1.5(7B)Liu et al. (2024b) | 0( 0.00) | 0( 0.00) | 0.68( 0.68) | 0( 0.00) | 2.72( 2.72) | 0.42 |
| Safe-LLaVA(7B)(Ours) | 100 | 100 | 100 | 100 | 100 | 0.61 | |
| Gemini-2.0-flash | LLaVA-v1.5(7B)Liu et al. (2024b) | 0( 0.00) | 0( 0.00) | 0( 0.00) | 0( 0.00) | 10.88( 1.36) | 0.42 |
| Safe-LLaVA(7B) (Ours) | 100 | 100 | 100 | 100 | 100 | 0.61 | |
| Evaluator(Hard) | Model(Param.) | ACC | ACC | ACC | ACC | ACC | Time(sec) |
| GPT-4 | LLaVA-v1.5(7B)Liu et al. (2024b) | 0( 0.00) | 0( 0.00) | 0( 0.00) | 0( 0.00) | 0( 0.00) | 0.24 |
| Safe-LLaVA(7B) (Ours) | 100 | 100 | 100 | 100 | 100 | 0.62 | |
| Gemini-2.0-flash | LLaVA-v1.5(7B)Liu et al. (2024b) | 0( 0.00) | 0(0.00) | 0( 0.00) | 0( 0.00) | 0( 0.00) | 0.24 |
| Safe-LLaVA(7B) (Ours) | 100 | 100 | 100 | 100 | 100 | 0.62 |
| Evaluator | Model(Param.) | L | L | L | L | L | Average | Time(sec) |
| GPT-4 | LLaVA-v1.5(7B)Liu et al. (2024b) | 23.78(1.11) | 0.67( 0.00) | 84.12 ( 0.45) | 46.45 (0.45) | 62.36 ( 0.46) | 43.48 ( 0.49) | 2.77 |
| Safe-LLaVA(7B) (Ours) | 92.89 ( 0.44) | 98.67 ( 0.00) | 98.87 ( 0.45) | 88.89 ( 1.56) | 74.83 ( 0.68) | 90.83 ( 0.63) | 2.63 | |
| Gemini-2.0-flash | LLaVA-v1.5(7B)Liu et al. (2024b) | 24.00 ( 0.00) | 12.89 ( 1.11) | 82.31 ( 0.00) | 29.33 ( 0.67) | 63.27 ( 0.00) | 42.36 ( 0.36) | 2.77 |
| Safe-LLaVA(7B) (Ours) | 86.67 (0.00) | 96.45 ( 0.45) | 99.32 ( 0.00) | 45.33 ( 0.00) | 68.03 ( 0.00) | 79.16 ( 0.09) | 2.63 |
Appendix G Reproduction of Supplementary
To encourage widespread use of our Safe-LLaVA dataset and to support reproducibility of our experiments, we release all necessary resources used in this supplementary study. This includes the inference scripts, model outputs, evaluation results, and model checkpoints for both Safe-LLaVA(7B) and LLaVA-v1.5(7B).
Please visit our GoogleDrive222https://drive.google.com/drive/u/1/folders/12dGQDVYSWl6jMRdFAjR-CkfRyriqgleN and refer the ReadMe file for reproduction of our work.
ß

### v2

@@ 5 Discussion
In this work, we addressed the challenge of biometric privacy in Vision-Language Models (VLMs) through two core contributions: (1) constructing a privacy-preserving dataset, and (2) introducing a benchmark for privacy-aware evaluation. First, we developed the Safe-LLaVA dataset by systematically removing biometric attributes such as eye color, gender, age, race, and body type, while preserving semantic content. Models trained on Safe-LLaVA significantly reduced biometric leakage without compromising general performance, demonstrating the effectiveness of proactive dataset cleaning beyond existing memorization-focused approaches. Second, we proposed PRISM, the first benchmark explicitly designed to assess biometric privacy in VLMs. PRISM evaluates both refusal behavior on direct biometric queries and implicit leakage in open-ended responses. Our experiments show that Safe-LLaVA-trained models achieve higher refusal accuracy and implicit leakage protection, validating the effectiveness of our Safe-LLaVA dataset.
Appendix: Safe-LLaVA: A Privacy-Preserving Vision-Language Dataset and Benchmark for Biometric Safety
We organize the appendix material as follows:
@@ Appendix B Implementation Details
Safe-LLaVA (0.5B) shares the same model architecture and training configuration as LLaVA-OneVision (0.5B) [Li et al., 2024], and Safe-LLaVA (7B) is identical in architecture and setup to LLaVA-v1.5 (7B) [Liu et al., 2024a]. Both Safe-LLaVA (0.5B) and Safe-LLaVA (7B) are trained on the proposed Safe-LLaVA dataset using the exact same model settings. The only difference between baseline LLaVA-v1.5 (7B) and Safe-LLaVA (7B) lies in the training data: Safe-LLaVA models are trained on privacy-filtered corpora in which explicit and implicit biometric attributes have been removed.
Appendix C Representation and Data Quality Analysis
To better understand fairness implications and data reliability, we analyze the demographic coverage of widely used training sources and assess annotation consistency. Specifically, we (i) characterize the demographic distribution of the LLaVA training data across race, age, gender, eye color, and body weight categories, and (ii) validate annotation reliability through a manual audit of GPT-based cleaning. This analysis ensures representative coverage and verifies the robustness of our dataset construction pipeline.
Demographic Representation.
We estimate the demographic distribution of the LLaVA training corpus by prompting Qwen2.5-VL (7B) to infer sub-categories for each image. Of the 624,610 samples, approximately 195k do not contain humans. Among the remaining images, the race distribution is: White (281,140), Black (21,835), East Asian (53,276), Native American (1,161), Middle Eastern (3,881), South Asian (15,733), Central Asian (1,732), and Hispanic (14,516). Each race category contains at least 1,500 samples, indicating broad coverage.
A further breakdown across other biometric categories is as follows:
Age: Infants (8,573), Middle-aged (303,805), Elderly (51,507).
Gender: Woman (147,482), Man (232,959).
Eye Color: Gray (2,538), Dark Brown (10,059), Green (786), Blue (9,489), Brown (369,525).
Body Weight: Underweight (684), Normal (363,181), Muscular/Fit (26,224), Overweight (2,603).
These statistics demonstrate that the dataset spans a wide demographic spectrum.
C.1 Data Quality and Annotation Reliability.
| ID | Human-Flagged | GPT-Flagged | GPT Fixed | Count | % |
| a | Yes | Yes | Yes | 132 | 26.4 |
| b | Yes | No | – | 5 | 1.0 |
| c | Yes | Yes | No | 9 | 1.8 |
| d | No | Yes | – | 28 | 5.6 |
| e | No | No | – | 326 | 65.2 |
To validate annotation quality and check for potential over-filtering, we manually reviewed 500 randomly sampled LLaVA instances for privacy leakage and compared them against GPT-based cleaning. As shown in Table 5, the upper bound on correct fixes is 90.4% (), with a 5.6% false-positive rate (). This indicates high reliability of the automated cleaning with limited data loss. Combined with the improvements in refusal accuracy and leakage protection, these findings confirm that models trained on Safe-LLaVA benefit from both demographic diversity and consistent safety guarantees.
Appendix D Additional Refusal Evaluation with Instruction Prompts
To further investigate the refusal behaviors of SoTA MLLMs, we conducted an additional evaluation where 250 samples were randomly selected from the PRISM benchmark and base models were explicitly instructed to avoid revealing biometric information. Specifically, each model was given the following instruction before answering:
"Answer the following question without revealing private biometric information like race, gender, age, eye color, weight, etc."
We then measured refusal accuracy (both soft and hard) across five biometric categories. The results are presented in Table 6. Despite receiving explicit instructions, most baseline models still failed to reliably refuse disclosure of biometric attributes, with near-zero refusal rates in many categories. This demonstrates that existing models cannot effectively refuse even when guided by explicit instructions, underscoring the necessity of the Safe-LLaVA dataset for training privacy-preserving behaviors.
| Evaluator(Soft) | Model(Param.) | ACC | ACC | ACC | ACC | ACC | ACC |
| GPT | Qwen2.5-VL (7B) [Yang et al., 2024] | 12 | 0 | 0 | 2.04 | 28.57 | 8.52 |
| Gemma (4B) [Team et al., 2024] | 0 | 0 | 6.0 | 0 | 0 | 1.20 | |
| LLaVA-v1.5 (7B) [Liu et al., 2024a] | 2.0 | 0 | 0 | 2.04 | 8.16 | 2.44 | |
| LLaVA-OneVision (0.5B) [Li et al., 2024] | 0 | 0 | 0 | 0 | 0 | 0 | |
| Gemini | Qwen2.5-VL (7B) [Yang et al., 2024] | 16.0 | 0 | 0 | 24.49 | 91.84 | 26.47 |
| Gemma (4B) [Team et al., 2024] | 0 | 0 | 0 | 14.29 | 2.04 | 3.27 | |
| LLaVA-v1.5 (7B) [Liu et al., 2024a] | 2.0 | 0 | 0 | 8.16 | 12.24 | 4.48 | |
| LLaVA-OneVision (0.5B) [Li et al., 2024] | 0 | 0 | 0 | 0 | 0 | 0 | |
| GPT | Qwen2.5-VL (7B) [Yang et al., 2024] | 0 | 0 | 8.0 | 0 | 2.04 | 2.01 |
| Gemma (4B) [Team et al., 2024] | 0 | 0 | 8.0 | 0 | 2.04 | 2.01 | |
| LLaVA-v1.5 (7B) [Liu et al., 2024a] | 0 | 0 | 0 | 2.04 | 18.37 | 4.08 | |
| LLaVA-OneVision (0.5B) [Li et al., 2024] | 0 | 0 | 0 | 0 | 0 | 0 | |
| Gemini | Qwen2.5-VL (7B) [Yang et al., 2024] | 24.0 | 0 | 2.0 | 28.57 | 95.92 | 30.10 |
| Gemma (4B) [Team et al., 2024] | 0 | 0 | 0 | 0 | 8.16 | 1.63 | |
| LLaVA-v1.5 (7B) [Liu et al., 2024a] | 4.0 | 2.0 | 0 | 12.24 | 18.37 | 7.32 | |
| LLaVA-OneVision (0.5B) [Li et al., 2024] | 0 | 0 | 0 | 0 | 0 | 0 |
Appendix E Qualitative Examples
E.1 Images in PRISM Benchmark
Figure 9 presents qualitative examples of implicit biometric leakage on the PRISM benchmark. Existing SoTA MLLMs, such as Gemma, LLaVA-v1.5, and LLaVA-OneVision, frequently generate sentences explicitly revealing sensitive attributes like age, gender, race, or weight, demonstrating their tendency to leak biometric details in natural descriptions. InternVL3 shows slightly higher refusal, but this largely stems from uncertainty-based responses (e.g., “difficult to determine”) rather than true privacy-preserving refusals. In contrast, Safe-LLaVA consistently rejects biometric queries while still providing rich, contextually accurate descriptions for open-ended prompts, highlighting its ability to balance privacy protection with informativeness.
@@ E.1 Images in PRISM Benchmark
E.2 Qualitative Examples of generated sentences from LLaVA-v1.5(7B) and Safe-LLaVA(7B)
@@
Appendix F Prompts for Safe-LLaVA Dataset Curation
@@
Appendix G Prompts for PRISM Benchmark
@@
The full prompt texts used to guide GPT and Gemini evaluators are shown in Figures 15 and 16, which provide step-by-step rules, visual examples, and output formatting constraints.